What are the intrinsic values and time premiums of the following call options if the price of the underlying stock is $35? What are the profits and losses to the buyers and the writers if the stock sells for $31 at the options' expiration?

Strike Price

Price of the Option

$30

$7.50

$35

$3.00
Answer #1

a.

Intrinsic Value of Call Option = Max(0, (Stock Price - Strike Price))

Intrinsic Value = (35 - 30)

Intrinsic value = $5

Time Value = Premium - Intrinsic Value

Time Value = 7.50 - 5

Time Value = $2.50

b.

Intrinsic Value = (35 - 35)

Intrinsic Value = 0

Time Value = 3 - 0

Time Value = $0